OpenStudy (anonymous):

According to the commutative property 7x-10y is equal to A. 10y-7x B.-10y+7x C.10x-7y D.-3xy

OpenStudy (anonymous):

commutative property refers to the order....

OpenStudy (anonymous):

Commutative Property states that a + b = b + a. In this case, 7x - 10 = -10 + 7x